What is the energy stored in the 17.5 ?F capacitor of a heart defibrillator charged to 8.25
Answer #1

a)

E = 1/2 *C*V^2 = 0.5*17.5*10^-6*8250 = 595.55 J

b)

Q = C*V = 17.5*10^-6*8250 = 0.1444 C = 144.4 mC